OpenStudy (anonymous):

Find the product of (2x + 5y)(2x − 5y).

ganeshie8 (ganeshie8):

use this :- \(\large (a+b)(a-b) = a^2 - b^2\)

OpenStudy (bahrom7893):

a = 2x, b = 5y

ganeshie8 (ganeshie8):

\( (2x + 5y)(2x − 5y) = (2x)^2 - (5y)^2 \) = ?

OpenStudy (anonymous):

4x2 − 25y2

OpenStudy (anonymous):

i say distibute

ganeshie8 (ganeshie8):

thats right ! gw !!

OpenStudy (anonymous):

4x^2-10xy+10xy-25y^2 4x^2-25y^2 so yes :)

OpenStudy (anonymous):

thank you :)

ganeshie8 (ganeshie8):

distribution always works :) yw !